OpenStudy (anonymous):

how about M-4/3 - 3M -1/5 =1 THE / IS A DIVISON SIGN ... SO ITS M-4 THEN UNDER THAT ITS THE 3

OpenStudy (anonymous):

-2M = 1+1/5+4/3 -2M = 38/15 M = -19/15

OpenStudy (anonymous):

m=7

OpenStudy (anonymous):

ah.. the old conundrum. pyt,baker please type out your question clearly.

OpenStudy (anonymous):

\[13+15=4m \]\[m-4/3 -3m-1/5= 1 \] like the denomenator \[5m-4-9m-9/15 =1\] \[-4m-13/15 =1\] \[13+15=4m\] 28=4m m=7

OpenStudy (anonymous):

SORRY DHATRADITYA I DO NOT KNOW HOW TO MAKE FACTION ON THIS WEB SITE THIS IS MY FIRST TIME

OpenStudy (anonymous):

the first line is mistaken the from 3rd line answer starts hope i could help!!

OpenStudy (anonymous):

thats alright, there is an equation button below on the left of the post button where you can type out your equations.

OpenStudy (anonymous):

THANK YOU I WILL TRY IT NEXT TIME IM SORRY HAVE TO GO NOW ... BUT THANK FOR ALL THE HELP

OpenStudy (anonymous):

hey i answerd ur question u thnx from him !!!!

OpenStudy (anonymous):

lol Roh, I thank you on her behalf :) have a medal

OpenStudy (anonymous):

lolz thnx ;)